OpenStudy (anonymous):

Show that the following matrix has only one distinct eigenvalue. What is it? (matrix in comments)

OpenStudy (anonymous):

\[C=\left[\begin{matrix}-1 & 0&1 \\ -1 & 0&1\\-1 & 0&1\end{matrix}\right]\]

OpenStudy (anonymous):

let me see if i we can work together as i've never done this yet i'm in linear algebra as of now

OpenStudy (anonymous):

Great - always good to have others working on the same material!

jimthompson5910 (jim_thompson5910):

\[\text{hint: the eigenvalues} \ \lambda \ \text{satisfy the equation:}\] \[\det\left(A-\lambda I\right)=0\]

OpenStudy (anonymous):

my books has the equatoion \[(\lambda I - A)x = 0\]

OpenStudy (anonymous):

as i haven't done determinants yet and the last time i did them was 4 years ago i might just tick with this equation

jimthompson5910 (jim_thompson5910):

hmm never seen it in that form

jimthompson5910 (jim_thompson5910):

that might be an intermediate step

OpenStudy (anonymous):

alright so i'll do it your way jim if i understand determinants right

OpenStudy (anonymous):

gahd i hate the equation editor how do you make a 3x3

OpenStudy (anonymous):

\[\det(A-\lambda I) = (-1-\lambda)(0-\lambda)(1-\lambda)+0+0-(1)(0-\lambda)(-1)-0-0=0\]

jimthompson5910 (jim_thompson5910):

Fill in the blanks \[\left[\begin{matrix}? & ? & ?\\ ? & ? & ?\\? & ? & ?\end{matrix}\right]\]

OpenStudy (anonymous):

Is that along the right lines? (sorry this is a bit confusing, helping both of us!)

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

looks right sorta

OpenStudy (anonymous):

that looks right

OpenStudy (anonymous):

if \[A=\left[\begin{matrix}a & b&c \\ d&e&f\\g&h&i\end{matrix}\right]\] then det(A) = (aei + bfg + cdh - ceg - bdi - afh) ... according to my notes!

OpenStudy (zarkon):

the characteristic polynomial is just \[-\lambda^3\] thus the only eigenvalue is 0

jimthompson5910 (jim_thompson5910):

When you take the determinant of \[\large A-\lambda I\], you should get \[\large -\lambda^3\]

OpenStudy (anonymous):

my book shows expanding the matrix by the first two colums

OpenStudy (anonymous):

which is what you have so if you correctly wrote it down, then yes that is correct

jimthompson5910 (jim_thompson5910):

yes, if feldy expands and simplifies then it'll come to negative lambda cubed

OpenStudy (anonymous):

Yep :) thanks!

OpenStudy (anonymous):

... OK if the next part of the question is "Obtain a complete set of independent eigenvectors of C. Your answer should be a set of column vectors." ... How would one go about answering that?

OpenStudy (anonymous):

Jim , got a calculator questiooon go on group chat

jimthompson5910 (jim_thompson5910):

ok one sec

jimthompson5910 (jim_thompson5910):

the eigenvectors that correspond to \[\large \lambda\] will be the vectors x that satisfy the equation \[\large A\vec{x}=\lambda \vec{x}\] But \[\large \lambda = 0\], so \[\large A\vec{x}=(0) \vec{x}\] which means that \[\large A\vec{x}=0\] So solve this equation to find your eigenvector

OpenStudy (anonymous):

Great, thanks so much!

OpenStudy (anonymous):

Sorry - have been busy between when this was first posted, and now.... Just getting back to this question... By using the Ax=0 equation, does that give you:\[\left[\begin{matrix}-1 & 0 &1\\-1 & 0 &1\\-1 & 0 &1 \end{matrix}\right] \left[\begin{matrix}x_1 \\ x_2 \\x_3\end{matrix}\right] = 0\] \[ie, -x_1+x_3 = 0\]\[x_1 = x_3\] .... How does this help in finding the eigenvectors??

OpenStudy (zarkon):

Looks like you have two free variables let \[x_3=t\text{ and }x_2=r\] write out the solution in terms of \[t\text{ and }r\]

OpenStudy (anonymous):

So my answer to this would be {[t,r,t]}??

OpenStudy (zarkon):

no...have you ever found the null space of a matrix?

OpenStudy (anonymous):

y...es? What do I do with it, once I found the null space? Thanks for the helpl (as always!)

OpenStudy (zarkon):

this problem turned into finding the null space \[A\vec{x}=0\]